OpenStudy (xochital):

What is the product of 7 the square root of 10. • 3the square root of 5? Simplify if possible

OpenStudy (anonymous):

7√(10) * 3√5 = 21√50 = 21√(25 * 2) = 21 * 5√2 = 105√2

OpenStudy (xochital):

thank you so much!

OpenStudy (anonymous):

Yeah (: Anything else

OpenStudy (xochital):

What is the simplified form of the expression \[\sqrt{4c ^{16}}-\sqrt{c ^{2}}+6\sqrt{c ^{2}}+\sqrt{9c}\]

OpenStudy (anonymous):

is their any multiple choices ?

OpenStudy (xochital):

oh yeah... \[2c ^{4} + 8c\]\[2c ^{4}+10c\]\[2c ^{8}+8c\]\[2c ^{8}+10c\]

OpenStudy (anonymous):

Okay I think it's b or d let me ask someone to help you okay . Dont want to give you the wrong answer

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

divide all the c exponents by 2 to undo the radical; then the simplification should be simpler

OpenStudy (amistre64):

and if those are the choices, then youve dropped an exponent in the post

OpenStudy (xochital):

oh yeah, the last one is supposed to be \[\sqrt{9c ^{2}}\]